2. A Rankine cycle has an initial mass of liquid flowing at 10 kg at 75℃. If final enthalpy is 379.6 KJ, find the change of the system.


1
Expert's answer
2022-03-12T02:02:15-0500

For pt1 t=75℃ h1= 313.93 kJ/kg

for pt2 h2 = 379.6 kJ/kg

∆h = (h2 − h1) = (379.6 - 313.93) = 65.67 kJ/kg

∆H = 65.67 kJ/kg × 10 kg = 656.7 kJ